OpenStudy (anonymous):

If a regular polygon has 54 diagonals how many sides does it have? What are the sums of the interior angles of the polygons? What are the sums of the exterior angles of the polygons? What is the measure of one interior angle? What is the measure of one exterior angle?

OpenStudy (anonymous):

the number of diagonals a polygon with s sides has is s C 2 - s = s*(s-1)/2 - s = s*(s-3)/2 12*(12-3)/2 = 54 12 sides

OpenStudy (anonymous):

Thanks, but what about the rest? I am not sure how to solve.